An eastbound bicyclist enters a tunnel at the same time a westbound bicyclist enters the other end of the tunnel. The eastbound bicyclist travels 10 km/hr and the westbound bicyclist travels 8 km/hr. A fly is flying back and forth between the two bicyclists at 15 km/hr. The fly leaves the eastbound bicyclist just as he enters the tunnel. The tunnel is 9 km long. How far has the fly traveled when the bicyclists meet in the tunnel?: An eastbound bicyclist enters a tunnel at the same time a westbound bicyclist enters the other end of the tunnel. The eastbound bicyclist travels 10 km/hr and the westbound bicyclist travels 8 km/hr. A fly is flying back and forth between the two bicyclists at 15 km/hr. The fly leaves the eastbound bicyclist just as he enters the tunnel. The tunnel is 9 km long. How far has the fly traveled when the bicyclists meet in the tunnel?
Answer by stanbon(19692)

An eastbound bicyclist enters a tunnel at the same time a westbound bicyclist enters the other end of the tunnel.
The eastbound bicyclist travels 10 km/hr and the westbound bicyclist travels 8 km/hr.
A fly is flying back and forth between the two bicyclists at 15 km/hr.
The fly leaves the eastbound bicyclist just as he enters the tunnel.
The tunnel is 9 km long. How far has the fly traveled when the bicyclists meet in the tunnel?
------------------
Find the time required before the cyclists meet:
Eastbound DATA:
rate = 10 km/h ; distance = x km ; time = x/10 hrs
-----------------
West bound DATA:
rate = 8 km/h ; distance = (9-x) km; time = (9-x)/8 hrs
-------------------
EQUATION:
time = time
x/10 = (9-x)/8
8x = 90-10x
18x = 90
x = 5 km
------------
Time before the two meet = x/10 = 5/10 = 1/2 hr.
---------------
In that time the fly will have flown (1/2)hr * 15/km/hr = 7.5 km
=======================================
Cheers
Stan H.